Teacher Li bought a total of 50 apples and oranges. If you buy 10 more oranges, the number of oranges will be twice that of apples. How many grass fruits and oranges did Teacher Li buy?


1
Expert's answer
2021-08-05T17:21:35-0400

Suppose teacher Li bought "x" apples. Then he bought "50-x" oranges.

If you buy 10 more oranges, the number of oranges will be twice that of apples


"2x=50-x+10"

"2x+x=60"

"3x=60"

"x=20"

"50-20=30"

Teacher Li bought "20" apples and "30" oranges.